A three-phase system has an effective line

voltage of 240 V, and is connected to a

balanced delta load which draws 5 kW at a

power factor of 0.8, then the line current is:


Select one:

a. 13.88 A

b. 16.4 A

c. 12.89 A

d. 15.04 A

e. 12.03 A


1
Expert's answer
2021-06-07T05:49:37-0400

Since 

"PF = \frac{P}{S} \implies S= \frac{ 5000}{0.8} - 6250 VA" .

From the formula for apparent power 

"S = 3U_pI_p \implies I_p=\frac{ 6250}{3*240}= 8.681A."  

Line current is "\sqrt{3}" times greater then phase current 

"I_L=\sqrt{3}*8.681 =15.035 A" (Option D)